OpenStudy (anonymous):

Identify the 19th term of the arithmetic sequence 5, 16, 27 Question 1 options: 1) 176 2) 214 3) 167 4) 203

OpenStudy (dumbcow):

203

OpenStudy (anonymous):

203

OpenStudy (angela210793):

16-5=11=d y19=y1+(n-1)*d=5+18*11=5+198=203